An exam consits of four true/false questions followed by five multiple choice questions each with 3 answers.
A) How many ways can a student answer the exam if they answer all of the questions?

B) How many ways can a student answer the exam if they can leave true/false questions blank?
C) How many ways can a student answer the exam if they can leave any of the questions blanks?


Answer by stanbon(57307)

An exam consits of four true/false questions followed by five multiple choice questions each with 3 answers.
----------------------------------------
A) How many ways can a student answer the exam if they answer all of the questions?
4 true/false in 2^4 = 16 ways
5 multiple choice in 3^5 = 243
Total: 16*243 = 3888
----------------------------------
B) How many ways can a student answer the exam if they can leave true/false questions blank?
4 true/false in 3^4 = 81
Total: 81*243 = 19683
--------------------------------------------
C) How many ways can a student answer the exam if they can leave any of the questions blanks?
Ans: 3^4*4^5
====================
cheers,
Stan H.